OpenStudy (anonymous):

yz has endpoints y(-1 -11) and z(-1 -3) find the length of yz to the nearest tenth

OpenStudy (anonymous):

@kohai @bibby @mathislife1234

OpenStudy (bibby):

distance formula = \(d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\)

OpenStudy (bibby):

plug in the coordinates. (x1,y1)=y(-1 -11) (x2,y2)=(-1 -3)

OpenStudy (anonymous):

would it be 12.36? or am i way off?

OpenStudy (bibby):

show your work

OpenStudy (anonymous):

am i wrong or right? i can show my answer if you tell me please

OpenStudy (bibby):

you're wrong

OpenStudy (anonymous):

aww okay then i dont wanna show my work if its wrong :/ can you help me out a little bit more please?

OpenStudy (bibby):

I don't think so. Just plug in given the information I gave you and you should get the answer

OpenStudy (anonymous):

1.41?

OpenStudy (anonymous):

my teacher is sitting right next to me and hes not helping me

OpenStudy (bibby):

@Analoha this problem is nothing but order of operations + plugging in you have the fact that \(\large d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) and the fact that x1= -1 ||| y1= -11 x2 = -1 ||| y2 = -3